Project

General

Profile

Actions

June 29th 2014

Goal : To do an activation run at 4V beam loss double rate
To do a TOF calibration
To do a CKOV HV scan and threshold scan

Issues encountered in setup and hall closure

  • Initial difficulty putting Q1 and Q2 into remote mode due to the operators not resetting the interlocks properly
    - Problem was corrected before search of the hall was completed
    - Magnet operators will need to be reminded periodically of magnet power on procedures.
  • The high voltage crates were not communicating with the high voltage control
    - Was traced to a flakey serial port connection. A more robust solution will have to be found.
  • TOF2 ADCs and KL TDCs were not available for run.
    - Was the result of partial progress in the reorganization the DAQ trigger.
    - Will be corrected with the completion of the work on the DAQ trigger.

Issues encountered during the run

  • Neutron monitor readout broken
    • webcam also pointing in the wrong direction (i.e. no backup in place for EPICS readout)
    • webcam fixed during a Controlled Entry
  • Confusion over shifter rostering
    • 1 left early
    • other missed lunch
  • Confusion regarding ISIS and MICE operations
    • ISIS rep rate and unexpected losses in SP 1 & 2
    • ISIS BLM app still unavailable via C&M menu
  • Run control not available
    - In progress upgrades to the MLCR computing systems had not filtered into run control.
    - Run control was not tested with DAQ because of Ckov HV problems (above).
    - Explicit instructions for bringing the run control up from an "off" or "fail" state not available.
  • DAQ produced fatal trigger mismatches on almost all runs.
    - Yordan suggested to cycle the power in the affected TDC crate and to remove and reinsert the affected Flash ADC card
    - Resulted in a long (5-10 minutes) turn around between runs because the procedure would not work every time.
    - Problem is believed to be hardware related; possibly in the TDC crate power supply. Component will need to be identified and fixed.
  • Q123 control crate wrongly turned off at end of run (significant chance of damaging power supply)
    - Instructions in MLCR. Shifters should consult documentation if in doubt.
  • Poor and confused preparation
    • MOM did not start Hall preparation work sufficiently in advance (~6 weeks)
    • Interim workarounds still not documented on Web_Whiteboard even since last October

Updated by Nebrensky, Henry over 9 years ago ยท 3 revisions